How do I find h & w when h is 7 more than w, and I am only given the area of 330ft^2?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



You are not "only given the area...", you are also given that is 7 more than .



You also know that the area is given by , so substitute:



And with the given value for area:



Put it in standard form:



Solve the quadratic (it factors, 22 and 15 have a difference of 7, hmmm), discard the negative root. The positive root is and is 7 more than that.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

How do I find h & w when h is 7 more than w, and I am only given the area of 330ft^2?
-------
Let width = w
Then height = (w + 7)
----
Area = width*height = 330 ft^2
-----
w^2+7w-330 = 0
---
Factor:
(w-15)(w+22) = 0
---
Width = 15 ft
Height = 15+7 = 22 ft
===========================
cheers,
Stan H.
==================